Question 13 What is the sound intensity level (in units of dB) when the intensity of sound is 5.78 mW/m2?
Answer #1

Sound intensity level is given by,

SL = 10*log (I/I0)

here,

I0 = threshold hearing level = 10^-12 W/m^2

I = Intensity of sound = 5.78 mW/m^2 = 5.78*10^-3 W/m^2

then,

SL = 10*log[(5.78*10^-3)/(10^-12)]

SL = 10*log(5.78*10^9) = 97.6193 dB

SL = 97.6 dB

"Let me know if you have any query."